# EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting

### v1

@@ A.4 Extended Results
Dataset T EntroPE iTrans. PatchTST TimesNet DLinear FEDformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.317 0.354 0.334 0.368 0.352 0.374 0.338 0.375 0.346 0.374 0.379 0.419 0.505 0.475
192 0.360 0.379 0.377 0.391 0.390 0.393 0.374 0.387 0.382 0.391 0.426 0.441 0.553 0.496
336 0.388 0.399 0.426 0.420 0.421 0.414 0.410 0.411 0.415 0.415 0.445 0.459 0.621 0.537
720 0.446 0.433 0.491 0.459 0.462 0.449 0.478 0.450 0.473 0.451 0.543 0.490 0.671 0.561
Avg. 0.378 0.391 0.407 0.410 0.406 0.407 0.400 0.406 0.404 0.408 0.448 0.452 0.588 0.517
ETTm2 96 0.180 0.265 0.180 0.264 0.183 0.270 0.187 0.267 0.193 0.293 0.203 0.287 0.255 0.339
192 0.241 0.311 0.250 0.309 0.255 0.314 0.249 0.309 0.284 0.361 0.269 0.328 0.281 0.340
336 0.313 0.352 0.311 0.348 0.309 0.347 0.321 0.351 0.382 0.429 0.325 0.366 0.339 0.372
720 0.411 0.413 0.412 0.407 0.412 0.404 0.408 0.403 0.558 0.525 0.421 0.415 0.433 0.432
Avg. 0.286 0.335 0.288 0.332 0.290 0.334 0.291 0.333 0.354 0.402 0.305 0.349 0.327 0.371
ETTh1 96 0.375 0.399 0.386 0.405 0.460 0.447 0.384 0.402 0.397 0.412 0.395 0.424 0.449 0.459
192 0.423 0.425 0.441 0.436 0.512 0.477 0.436 0.429 0.446 0.441 0.469 0.470 0.500 0.482
336 0.429 0.432 0.487 0.458 0.546 0.496 0.638 0.469 0.489 0.467 0.490 0.477 0.521 0.496
720 0.439 0.454 0.503 0.491 0.544 0.517 0.521 0.500 0.513 0.510 0.598 0.544 0.514 0.512
Avg. 0.416 0.425 0.454 0.447 0.516 0.484 0.495 0.450 0.461 0.457 0.498 0.484 0.496 0.487
ETTh2 96 0.281 0.336 0.297 0.349 0.308 0.355 0.340 0.374 0.340 0.394 0.358 0.397 0.346 0.388
192 0.371 0.393 0.380 0.400 0.393 0.405 0.402 0.414 0.482 0.479 0.429 0.439 0.456 0.452
336 0.392 0.394 0.428 0.432 0.427 0.436 0.452 0.452 0.591 0.541 0.496 0.487 0.482 0.486
720 0.421 0.427 0.427 0.445 0.436 0.450 0.462 0.468 0.839 0.661 0.463 0.474 0.515 0.511
Avg. 0.366 0.387 0.383 0.407 0.391 0.411 0.414 0.427 0.563 0.519 0.437 0.449 0.450 0.459
Weather 96 0.164 0.211 0.174 0.214 0.186 0.227 0.172 0.220 0.195 0.252 0.217 0.296 0.266 0.336
192 0.210 0.252 0.221 0.254 0.234 0.265 0.219 0.261 0.237 0.295 0.276 0.336 0.307 0.367
336 0.256 0.290 0.278 0.296 0.284 0.301 0.246 0.337 0.282 0.331 0.339 0.380 0.359 0.395
720 0.339 0.342 0.358 0.347 0.356 0.349 0.365 0.359 0.345 0.382 0.403 0.428 0.419 0.428
Avg. 0.242 0.273 0.258 0.278 0.265 0.285 0.251 0.294 0.265 0.315 0.309 0.360 0.338 0.382
Electricity 96 0.163 0.252 0.148 0.240 0.190 0.296 0.168 0.272 0.210 0.302 0.193 0.308 0.201 0.317
192 0.177 0.268 0.162 0.253 0.199 0.304 0.184 0.322 0.210 0.305 0.201 0.315 0.222 0.334
336 0.194 0.284 0.178 0.269 0.217 0.319 0.198 0.300 0.223 0.319 0.214 0.329 0.231 0.443
720 0.235 0.321 0.225 0.317 0.258 0.352 0.220 0.320 0.258 0.350 0.246 0.355 0.254 0.361
Avg. 0.182 0.271 0.178 0.270 0.216 0.318 0.193 0.304 0.225 0.319 0.214 0.327 0.227 0.338
Exchange 96 0.083 0.140 0.086 0.206 0.088 0.205 0.107 0.234 0.088 0.218 0.148 0.278 0.197 0.323
192 0.177 0.235 0.177 0.299 0.176 0.299 0.226 0.344 0.176 0.315 0.271 0.315 0.300 0.369
336 0.299 0.371 0.331 0.417 0.301 0.397 0.367 0.448 0.313 0.427 0.460 0.427 0.509 0.524
720 0.761 0.869 0.847 0.691 0.901 0.714 0.964 0.746 0.839 0.695 1.195 0.695 1.447 0.941
Avg. 0.331 0.386 0.360 0.403 0.367 0.404 0.416 0.443 0.354 0.414 0.519 0.429 0.613 0.539
A.4.2 FilterTS Performance Comparison Across Configurations
@@
Exchange Rate 96 96 0.085 0.201 0.083 0.200 0.084 0.201 0.082 0.198
96 192 0.180 0.300 0.171 0.294 0.173 0.295 0.172 0.294
96 336 0.342 0.420 0.323 0.411 0.316 0.407 0.322 0.410
96 720 0.878 0.699 0.916 0.713 0.827 0.685 0.828 0.686
Avg. 0.371 0.405 0.373 0.405 0.350 0.397 0.351 0.397
A.5 Entropy Qualitative Analysis
@@
A.6 Time Series Quantization
@@
A.7 Entropy Calculation
@@
A.8 Cross-Entropy Loss
@@
A.9 Mean Squared Error (MSE)
@@
A.10 EntroPE Model Algorithms
Algorithm 1 EntroPE Forward Pass
1:Raw time series $X\in\mathbb{R}^{B\times C\times L}$, forecast horizon $T$
2:Predictions $\hat{Y}\in\mathbb{R}^{B\times C\times T}$
[…]
14:return $\hat{Y}$
Algorithm 2 Entropy-Based Dynamic Patcher (EDP)
1:Tokenized sequence $X^{tok}\in\mathbb{R}^{(B\cdot C)\times L}$, pre-trained entropy model $\theta$
2:Patch boundaries $\mathcal{B}$, patch mask $M\in\mathbb{R}^{(B\cdot C)\times L}$
3:Initialize $\mathcal{B}\leftarrow\emptyset$, $M\leftarrow\mathbf{0}$
4:Phase 1: Entropy Calculation
5:for $t=1$ to L1L-1 do
6:  $\text{context}\leftarrow X^{tok}[:,1:t]$
7:  $\text{logits}\leftarrow\text{entropy\_model}_{\theta}(\text{context})$
8:  $p_{\theta}(x_{t+1}|x_{\leq t})\leftarrow\text{softmax}(\text{logits})$
9:  $H(x_{t})\leftarrow-\sum_{v\in\mathcal{V}}p_{\theta}(v)\log p_{\theta}(v)$
10:end for
11:Phase 2: Boundary Detection
12:for $t=2$ to L1L-1 do
13:  if $H(x_{t})>\theta$ and $H(x_{t})-H(x_{t-1})>\gamma$ then
14:   $\mathcal{B}\leftarrow\mathcal{B}\cup\{t\}$;  $M[:,t]\leftarrow 1$
15:  end if
16:end for
17:Add first and last positions as boundaries
[…]
19:M[:,1]1M[$\leftarrow 1$;  $M[:,L]\leftarrow 1$
20:return $\mathcal{B},M$
Algorithm 3 Adaptive Patch Encoder (APE)
1:Continuous sequence $X^{ci}\in\mathbb{R}^{(B\cdot C)\times L}$, patch boundaries $\mathcal{B}$, vocab size $V$
2:Patch embeddings $P^{in}$, hidden states $H^{enc}$
[…]
5:Extract variable-length patches
[…]
7:for $i=1$ to $|\mathcal{B}|-1$ do
8:  $\text{patch}_{i}\leftarrow X^{ci}[:,\mathcal{B}[i]:\mathcal{B}[i+1]]$
9:  Append $\text{patch}_{i}$ to patches
10:end for
11:Initial patch embedding via max pooling
12:for each patch $p_{j}$ do
13:  $\text{emb}\leftarrow E(\text{tokenize}(p_{j}))$
14:  $P_{0}[:,j,:]\leftarrow\text{MaxPool}(\text{emb})$
15:end for
16:Cross-attention refinement (NN layers)
[…]
18:for $n=1$ to NN do
19:  $h_{n}\leftarrow\text{TransformerEncoder}_{n}(h_{\text{curr}})$
20:  $Q\leftarrow W_{q}(P_{\text{curr}})$
21:  $K\leftarrow W_{k}(h_{n})$,   $V\leftarrow W_{v}(h_{n})$
22:  $\text{out}\leftarrow\text{softmax}\!\left(\tfrac{QK^{T}}{\sqrt{d_{k}}}\right)V$
23:  $P_{n}\leftarrow P_{\text{curr}}+W_{o}(\text{out})$
24:  Update $h_{\text{curr}},P_{\text{curr}}$
25:end for
26:return $P^{in}\leftarrow P_{N}$,   $H^{enc}\leftarrow h_{N}$
A.11 Use of Large Language Models

### v2

@@ E.6 Training Efficiency Analysis
Dataset T EntroPE TSLANet iTrans. PatchTST TimesNet DLinear FEDformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.317 0.354 0.321 0.362 0.334 0.368 0.352 0.374 0.338 0.375 0.346 0.374 0.379 0.419 0.505 0.475
192 0.360 0.379 0.361 0.383 0.377 0.391 0.390 0.393 0.374 0.387 0.382 0.391 0.426 0.441 0.553 0.496
336 0.388 0.399 0.382 0.404 0.426 0.420 0.421 0.414 0.410 0.411 0.415 0.415 0.445 0.459 0.621 0.537
720 0.446 0.433 0.446 0.438 0.491 0.459 0.462 0.449 0.478 0.450 0.473 0.451 0.543 0.490 0.671 0.561
Avg. 0.378 0.391 0.378 0.397 0.407 0.410 0.406 0.407 0.400 0.406 0.404 0.408 0.448 0.452 0.588 0.517
ETTm2 96 0.180 0.265 0.179 0.261 0.180 0.264 0.183 0.270 0.187 0.267 0.193 0.293 0.203 0.287 0.255 0.339
192 0.241 0.311 0.243 0.303 0.250 0.309 0.255 0.314 0.249 0.309 0.284 0.361 0.269 0.328 0.281 0.340
336 0.313 0.352 0.308 0.345 0.311 0.348 0.309 0.347 0.321 0.351 0.382 0.429 0.325 0.366 0.339 0.372
720 0.411 0.413 0.403 0.400 0.412 0.407 0.412 0.404 0.408 0.403 0.558 0.525 0.421 0.415 0.433 0.432
Avg. 0.286 0.335 0.283 0.327 0.288 0.332 0.290 0.334 0.291 0.333 0.354 0.402 0.305 0.349 0.327 0.371
ETTh1 96 0.375 0.399 0.387 0.405 0.386 0.405 0.460 0.447 0.384 0.402 0.397 0.412 0.395 0.424 0.449 0.459
192 0.423 0.425 0.448 0.436 0.441 0.436 0.512 0.477 0.436 0.429 0.446 0.441 0.469 0.470 0.500 0.482
336 0.429 0.432 0.451 0.437 0.487 0.458 0.546 0.496 0.638 0.469 0.489 0.467 0.490 0.477 0.521 0.496
720 0.439 0.454 0.505 0.485 0.503 0.491 0.544 0.517 0.521 0.500 0.513 0.510 0.598 0.544 0.514 0.512
Avg. 0.416 0.425 0.448 0.441 0.454 0.447 0.516 0.484 0.495 0.450 0.461 0.457 0.498 0.484 0.496 0.487
ETTh2 96 0.281 0.336 0.299 0.354 0.297 0.349 0.308 0.355 0.340 0.374 0.340 0.394 0.358 0.397 0.346 0.388
192 0.371 0.393 0.382 0.399 0.380 0.400 0.393 0.405 0.402 0.414 0.482 0.479 0.429 0.439 0.456 0.452
336 0.392 0.394 0.379 0.396 0.428 0.432 0.427 0.436 0.452 0.452 0.591 0.541 0.496 0.487 0.482 0.486
720 0.421 0.427 0.429 0.445 0.427 0.445 0.436 0.450 0.462 0.468 0.839 0.661 0.463 0.474 0.515 0.511
Avg. 0.366 0.387 0.372 0.399 0.383 0.407 0.391 0.411 0.414 0.427 0.563 0.519 0.437 0.449 0.450 0.459
Weather 96 0.164 0.211 0.177 0.217 0.174 0.214 0.186 0.227 0.172 0.220 0.195 0.252 0.217 0.296 0.266 0.336
192 0.210 0.252 0.224 0.257 0.221 0.254 0.234 0.265 0.219 0.261 0.237 0.295 0.276 0.336 0.307 0.367
336 0.256 0.290 0.278 0.297 0.278 0.296 0.284 0.301 0.246 0.337 0.282 0.331 0.339 0.380 0.359 0.395
720 0.339 0.342 0.355 0.347 0.358 0.347 0.356 0.349 0.365 0.359 0.345 0.382 0.403 0.428 0.419 0.428
Avg. 0.242 0.273 0.259 0.280 0.258 0.278 0.265 0.285 0.251 0.294 0.265 0.315 0.309 0.360 0.338 0.382
Electricity 96 0.163 0.252 0.175 0.260 0.148 0.240 0.190 0.296 0.168 0.272 0.210 0.302 0.193 0.308 0.201 0.317
192 0.177 0.268 0.182 0.268 0.162 0.253 0.199 0.304 0.184 0.322 0.210 0.305 0.201 0.315 0.222 0.334
336 0.194 0.284 0.199 0.285 0.178 0.269 0.217 0.319 0.198 0.300 0.223 0.319 0.214 0.329 0.231 0.443
720 0.235 0.321 0.240 0.317 0.225 0.317 0.258 0.352 0.220 0.320 0.258 0.350 0.246 0.355 0.254 0.361
Avg. 0.182 0.271 0.199 0.283 0.178 0.270 0.216 0.318 0.193 0.304 0.225 0.319 0.214 0.327 0.227 0.338
E.6.2 FilterTS Performance Comparison Across Configurations
@@
Appendix F Entropy Qualitative Analysis
@@
F.1 Time Series Quantization
@@
Automatic Determination of Quantization Range.
We avoid manually choosing a fixed tokenization range (e.g., [8,8][-8,8]) and instead determine the quantization interval automatically from the training data. This ensures that the discretization step is fully data-driven and does not require hand-tuning for each dataset.
$\gamma_{\text{rel}}=Q_{\alpha}\left({\Delta H(x_{t})}_{t=2}^{T}\right)$ (39)
First, all continuous features are z-score normalized using the mean and standard deviation computed from the training split only. Let $X_{\text{train}}$ denote the collection of all normalized values that are passed to the tokenizer. We compute two empirical quantiles:
$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\qquad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}}),$
where $\epsilon$ controls the desired coverage (we use $\epsilon=0.005$, corresponding to $99.5\%$ of the training distribution). We then define a symmetric quantization radius
$R=\max\left(|q_{\text{low}}|,\;|q_{\text{high}}|\right),$
and construct the discretization interval as [R,R][-R,R].
This interval is uniformly divided into $V$ bins (we use $V=256$ unless otherwise stated). Values falling outside [R,R][-R,R] are clipped to the nearest boundary bin. This procedure allows the tokenizer to adapt automatically to the empirical scale and variability of the dataset, providing a robust discretization scheme without manual hyperparameter selection.
Appendix G Utility & Loss Functions
G.1 Entropy Calculation
@@
G.2 Cross-Entropy Loss
@@
G.3 Mean Squared Error (MSE)
@@
Use of Large Language Models